-9x+12y=18

x-2y=3

Answer by rmromero(383)


-9x+12y=18
x-2y=3
Here are the Steps using Substition method
1. In either equation, solve for one variable in terms of the other.
x - 2y = 3
x = 2y + 3
2. Substitute for that variable in the other equation. Solve.
-9x + 12y = 18, x = 2y + 3
-9(2y + 3) + 12y = 18
-18y - 27 + 12y = 18
-6y = 45
y = -45/6
= -15/2
3. Substitute the result from step 2 in either equation. Solve for the other variable.

x - 2y = 3 , y = -15/2
x - 2(-15/2) = 3
x - (-15) = 3
x + 15 = 3
x = -12

4. Check the solution in both original equations.

x - 2y = 3, x = -12 and y = -15/2

-12 - 2(-15/2) = 3
-12 + 15 = 3
3 = 3 ------> true

-9x + 12y = 18, x = -12 and y = -15/2
-9(-12) + 12(-15/2) = 18
108 - 90 = 18
18 = 18 ----> True

Therefore, the solution is x = -12 and y = -15/2


Visit my website:
http://onlinemathtutorial.tripod.com
Thanks